OpenStudy (anonymous):

The word "independent" is spelled out on index cards, one letter per card. Ken randomly selects a card and does not replace it, and then Barb randomly selects a card. What is the probability that they both get a card with an "e" written on it? A. 3/55 B. 9/121 C. 26/55 D. 6/11 I got A!

Directrix (directrix):

There are 3 of the letter E in the word independent. There are 11 letters in the word "independent."

Directrix (directrix):

The probabilty that Ken draws the letter E = the number of Es divided by the number of letters. That would be 3/11. The probability that Barb draws the letter E = the number of remaining Es divided by the number of remaining letters. That would be 2/10. P(Ken E and Barb E) = 3/11 * 2/10 = ?

Directrix (directrix):

>>> I got A! I agree with that, @katlin95